# A Decentralized Pilot Assignment Algorithm for Scalable O-RAN Cell-Free Massive MIMO

### v4

@@ Abstract
Index Terms:
Open-RAN (O-RAN), cell-free massive MIMO, deep reinforcement learning, pilot assignment.
@@ II-A CFmMIMO Configuration in O-RAN Architecture
Our decentralized O-RAN CFmMIMO system is illustrated in Fig. 1(b). We consider $M$ single-antenna O-RUs and $U$ O-DUs collected in sets =$\mathcal{M}=\{1,2,\ldots,M\}$ and $\mathcal{U}=\{1,2,\ldots,U\}$, respectively. The O-RUs are randomly placed using uniform distribution across the area that is divided into $U$ disjoint regions for system decentralization, and each O-DU is deployed to one of the regions. Each O-RU is connected to one of the O-DUs in $\mathcal{U}$ via an open fronthaul (O-FH) connection such that O-RUs within each subdivided region are connected to the same O-DU. We define $\mathcal{M}^{\mathsf{DU}}_{u}\subseteq\mathcal{M}$ as the set of O-RUs connected to O-DU $u\in\mathcal{U}$. We assume inter-DU connections [36] to form RU clusters that are fully user-centric since the users can be served by RUs from different sets of $\mathcal{M}^{\mathsf{DU}}_{u}$. We assume ideal O-FH and inter-DU connections [31, 30].
@@
Next, we consider $K$ single-antenna users in a set $\mathcal{K}=\{1,2,\ldots,K\}$. For each user $k$, a user-centric RU cluster is formed such that only $M^{\mathsf{UE}}_{k}\ll M$ O-RUs are engaged to serve the user, where we define $\mathcal{M}^{\mathsf{UE}}_{k}\subset\mathcal{M}$ to be the set of O-RUs serving user $k\in\mathcal{K}$ (i.e., $M^{\mathsf{UE}}_{k}=|\mathcal{M}^{\mathsf{UE}}_{k}|$ where ||$\cdot$|| ⋅ | denotes the set cardinality). Each $\mathcal{M}^{\mathsf{UE}}_{k}$ is assumed to be selected and updated using a procedure independent from our PA (e.g., radio resource control (RRC) setup procedure [37]). We also define $\mathcal{K}^{\mathsf{RU}}_{m}\subset\mathcal{K}$ to be the set of users served by O-RU $m\in\mathcal{M}$.
@@ II-C Codebook-based Channel Estimation
To conduct channel estimation, each user $k\in\mathcal{K}$ transmits the assigned pilot $\bm{x}^{(i)}_{k}$. The signal vector (of length $T_{p}$) received by O-RU $m\in\mathcal{M}$ is then expressed as $\bm{y}^{(i)}_{m}=\mathbf{X}^{(i)}\bm{g}^{(i)}_{m}+\bm{w}^{(i)}_{m}=\sum_{k\in\mathcal{K}}g^{(i)}_{km}\bm{x}^{(i)}_{k}+\bm{w}^{(i)}_{m}$, where $\mathbf{X}^{(i)}=[\bm{x}^{(i)}_{1}\bm{x}^{(i)}_{2}\cdots\bm{x}^{(i)}_{K}]$ is the $T_{p}\times K$ pilot matrix and $\bm{g}^{(i)}_{m}=[g^{(i)}_{1m}\;g^{(i)}_{2m}\cdots g^{(i)}_{Km}]^{\top}$ is the channel vector (of length $K$) for O-RU $m$. Here, $\bm{w}^{(i)}_{m}\sim\mathcal{CN}(\mathbf{0},\sigma^{2}\mathbf{I}_{T_{p}})$ is the zero-mean complex Gaussian noise vector of length $T_{p}$ with covariance $\sigma^{2}\mathbf{I}_{T_{p}}$, where $\mathbf{I}_{n}$ is the $n\times n$ identity matrix.
@@
$\widehat{g}^{(i)}_{km}=(\bm{x}_{k}^{(i)})^{\text{H}}\bm{y}^{(i)}_{m}=\sum_{k^{\prime}\in\mathcal{K}}g^{(i)}_{k^{\prime}m}(\bm{x}_{k}^{(i)})^{\text{H}}\bm{x}^{(i)}_{k^{\prime}}+(\bm{x}_{k}^{(i)})^{\text{H}}\bm{w}^{(i)}_{m}=g^{(i)}_{km}+\sum_{\begin{subarray}{c}k^{\prime}\in\mathcal{K}\\ k^{\prime}\neq k\end{subarray}}g^{(i)}_{k^{\prime}m}(\bm{x}_{k}^{(i)})^{\text{H}}\bm{x}^{(i)}_{k^{\prime}}+(\bm{x}_{k}^{(i)})^{\text{H}}\bm{w}^{(i)}_{m}.$ (4)
@@ II-D Data Transmission Model
$\mathsf{SINR}^{\mathsf{u}}_{k}=\frac{\tiny\rho_{k}\left|\mathbb{E}\left[\sum_{m\in{\mathcal{M}^{\mathsf{UE}}_{k}}}(\widehat{g}^{(i)}_{km})^{*}g^{(i)}_{km}\right]\right|^{2}}{\tiny\displaystyle\sum_{k^{\prime}\in\mathcal{K}}\rho_{k^{\prime}}\mathbb{E}\Bigg{[}\bigg{|}\sum_{m\in\mathcal{M}^{\mathsf{UE}}_{k}}(\widehat{g}^{(i)}_{km})^{*}g^{(i)}_{k^{\prime}m}\bigg{|}^{2}\Bigg{]}-\rho_{k}\Bigg{|}\mathbb{E}\Bigg{[}\sum_{m\in{\mathcal{M}^{\mathsf{UE}}_{k}}}(\widehat{g}^{(i)}_{km})^{*}g^{(i)}_{km}\Bigg{]}\Bigg{|}^{2}+\sigma^{2}_{\mathsf{u}}\sum_{m\in{\mathcal{M}^{\mathsf{UE}}_{k}}}\mathbb{E}\left[|\widehat{g}^{(i)}_{km}|^{2}\right]},$ (6)
@@
For downlink transmission, the data signal $x^{\mathsf{d}}_{k}$ is transmitted by the O-RUs serving user $k$ (i.e., O-RU $m\in\mathcal{M}^{\mathsf{UE}}_{k}$) after applying the conjugate beamforming expressed as $\bar{x}^{\mathsf{d}}_{km}=(\widehat{g}^{(i)}_{km})^{*}x^{\mathsf{d}}_{k}/|\widehat{g}^{(i)}_{km}|$. The received signal $\bar{y}^{\mathsf{d}}_{k}$ for user $k$ is then given by
@@ II-E Problem Formulation
$\mathsf{MSE}^{(i)}_{k}$ $=\mathbb{E}\bigg{[}\sum_{m\in\mathcal{M}^{\mathsf{UE}}_{k}}\Big{|}\widehat{g}^{(i)}_{km}-g^{(i)}_{km}\Big{|}^{2}\bigg{]}=\sum_{m\in\mathcal{M}^{\mathsf{UE}}_{k}}\mathbb{E}\left[\Big{|}\widehat{g}^{(i)}_{km}-g^{(i)}_{km}\Big{|}^{2}\right]$
$=\sum_{m\in\mathcal{M}^{\mathsf{UE}}_{k}}\mathbb{E}\Bigg{[}\Big{|}\sum_{\begin{subarray}{c}k^{\prime}\in\mathcal{K}\\ k^{\prime}\neq k\end{subarray}}g^{(i)}_{k^{\prime}m}(\bm{x}_{k}^{(i)})^{\text{H}}\bm{x}^{(i)}_{k^{\prime}}+(\bm{x}_{k}^{(i)})^{\text{H}}\bm{w}^{(i)}_{m}\Big{|}^{2}\Bigg{]}=\sum_{m\in\mathcal{M}^{\mathsf{UE}}_{k}}\sum_{\begin{subarray}{c}k^{\prime}\in\mathcal{K}\\ k^{\prime}\neq k\end{subarray}}\beta^{(i)}_{k^{\prime}m}\Big{|}(\bm{x}_{k}^{(i)})^{\text{H}}\bm{x}^{(i)}_{k^{\prime}}\Big{|}^{2}+\sigma^{2},$ (9)
@@
[…]
$\hskip 17.07164pt\|\bm{\phi}^{(i)}_{u,t}\|^{2}_{2}=1,\left(\bm{\phi}_{u,t}^{(i)}\right)^{\text{H}}\bm{\phi}^{(i)}_{u,t^{\prime}}=0\text{ if }t\neq t^{\prime},\;\;\forall t,t^{\prime}=1,2,\ldots,T_{p}.$ (12)
@@ III-B Preliminaries on Multi-agent Deep Reinforcement Learning
MA-DRL addresses scenarios where multiple agents perform simultaneous decision-making based on a Markov game model [44]. For our decentralized PA problem, we define MA-DRL using a tuple ($(\{\bm{S}^{(\ell)}_{u}\}_{u\in\mathcal{U}}$, $\{\bm{a}^{(\ell)}_{u}\}_{u\in\mathcal{U}}$, $\{r^{(\ell)}_{u}\}_{u\in\mathcal{U}})$, where $\bm{S}^{(\ell)}_{u}$, $\bm{a}^{(\ell)}_{u}$, and $r^{(\ell)}_{u}$ are respectively the state, action, and reward of the agent $u$ during the $\ell$-th near-RT loop. For each loop $\ell$, agent $u$ with a state $\bm{S}^{(\ell)}_{u}$ makes an action $\bm{a}^{(\ell)}_{u}$ to interact with the environment. Subsequently, the agent makes an observation and computes a reward $r^{(\ell)}_{u}$ which helps to find the next state $\bm{S}^{(\ell+1)}_{u}$.
In the non-RT loop, once an agent has completed multiple interactions with the environment, its policy on action selection for a given state is optimized by updating the weights of its respective deep neural network (DNN). Here the action is determined based on the Q-value [45] denoted by $Q(\bm{S}^{(\ell)}_{u},\bm{a}^{(\ell)}_{u})$. The Q-value quantifies the quality of an agent’s action for a given state. Thus, it is important for the agent to obtain accurate Q-values to make correct decisions. In DRL, these Q-values are computed via a DNN, the weights of which are trained with experiences so that a correct (i.e., Q-value-maximizing) action can be selected upon each decision-making.
@@ III-C1 States
$[\bm{\Phi}_{u}^{(\ell)}]_{k,t}=\begin{cases}1&\text{if }\bm{x}^{(N_{\mathsf{n}}\ell)}_{k}=\bm{\phi}^{(N_{\mathsf{n}}\ell)}_{u,t},\\ 0&\text{otherwise}.\end{cases}$ (13)
@@ III-C3 Rewards
where $p=\widetilde{p}^{(\ell)}_{u}$ by the availability of inter-DU message. The mapping function (16) converts the observation data into a reward range such that lower values of $p$ are rewarded higher. Here $[p_{\text{min}},p_{\text{max}}]$ is the range of observation data, which we assume is set by the non-RT RIC.
@@
First, in terms of the parameters defined in our model, we find the expected reward at near-RT loop $\ell$ for a given state-action pair $(\bm{S}^{(\ell)}_{u},\bm{a}^{(\ell)}_{u})$, which is expressed as
@@
$\bm{a}_{u}^{(\ell)}=\operatorname*{arg\,min}_{\bm{a}_{u}\in\mathcal{A}_{u}}\;\sum_{k\in\mathcal{K}^{\mathsf{DU}}_{u}}\sum_{m\in\mathcal{M}^{\mathsf{UE}}_{k}}\mathbb{E}[p^{(\ell)}_{km}]=\operatorname*{arg\,min}_{\bm{a}_{u}\in\mathcal{A}_{u}}\frac{1}{N_{\mathsf{n}}}\sum_{k\in\mathcal{K}^{\mathsf{DU}}_{u}}\sum_{m\in\mathcal{M}^{\mathsf{UE}}_{k}}\sum_{n=0}^{N_{\mathsf{n}}-1}\mathbb{E}\left[\left|\widehat{g}^{(N_{\mathsf{n}}\ell+n)}_{km}\right|^{2}\right],$ (21)
@@ III-D MA-DRL-based Pilot Assignment Scheme
1 Input: Pilot length $T_{p}$, number of RT loops $N$, number of RT loops per near-RT loop $N_{\mathsf{n}}$, number of internal loops $L$, set of users managed by O-DU $u$ $\mathcal{K}^{\mathsf{DU}}_{u}$, set of O-RUs managed by O-DU $u$ $\mathcal{M}^{\mathsf{DU}}_{u}$, set of users served by O-RU $m$ $\mathcal{K}^{\mathsf{RU}}_{m}$, set of O-RUs serving the user $k$ $\mathcal{M}^{\mathsf{UE}}_{k}$, training period, update period
@@ 13       Update the PA according to $\bm{a}^{(\ell)}_{u}$
16      if mod(,training $\text{mod}(\ell,\text{training period})=0$ then generate a batch from the memory and train $\bm{\theta}^{\mathsf{tr}}_{u}$ via SGD on (24)
17       if mod(,update $\text{mod}(\ell,\text{update period})=0$ then set $\bm{\theta}^{\mathsf{ta}}_{u}$ = $\bm{\theta}^{\mathsf{tr}}_{u}$

### v5

@@ II-A CFmMIMO Configuration in O-RAN Architecture
Our decentralized O-RAN CFmMIMO system is illustrated in Fig. 1(b). We consider $M$ single-antenna O-RUs and $U$ O-DUs collected in sets $\mathcal{M}=\{1,2,\ldots,M\}$ and $\mathcal{U}=\{1,2,\ldots,U\}$, respectively. The O-RUs are randomly placed using uniform distribution across the area that is divided into $U$ disjoint regions for system decentralization, and each O-DU is deployed to one of the regions. Each O-RU is connected to one of the O-DUs in $\mathcal{U}$ via an open fronthaul (O-FH) connection such that O-RUs within each subdivided region are connected to the same O-DU. We define $\mathcal{M}^{\mathsf{DU}}_{u}\subseteq\mathcal{M}$ as the set of O-RUs connected to O-DU $u\in\mathcal{U}$. We assume inter-DU connections [36] to form RU clusters that are fully user-centric since the users can be served by RUs from different sets of $\mathcal{M}^{\mathsf{DU}}_{u}$. We assume ideal O-FH and inter-DU connections [31, 30].
@@
Next, we consider $K$ single-antenna users in a set $\mathcal{K}=\{1,2,\ldots,K\}$. For each user $k$, a user-centric RU cluster is formed such that only $M^{\mathsf{UE}}_{k}\ll M$ O-RUs are engaged to serve the user, where we define $\mathcal{M}^{\mathsf{UE}}_{k}\subset\mathcal{M}$ to be the set of O-RUs serving user $k\in\mathcal{K}$ (i.e., $M^{\mathsf{UE}}_{k}=|\mathcal{M}^{\mathsf{UE}}_{k}|$ where $|\cdot|$ denotes the set cardinality). Each $\mathcal{M}^{\mathsf{UE}}_{k}$ is assumed to be selected and updated using a procedure independent from our PA (e.g., radio resource control (RRC) setup procedure [37]). We also define $\mathcal{K}^{\mathsf{RU}}_{m}\subset\mathcal{K}$ to be the set of users served by O-RU $m\in\mathcal{M}$.
@@ II-C Codebook-based Channel Estimation
To conduct channel estimation, each user $k\in\mathcal{K}$ transmits the assigned pilot $\bm{x}^{(i)}_{k}$. The signal vector (of length $T_{p}$) received by O-RU $m\in\mathcal{M}$ is then expressed as $\bm{y}^{(i)}_{m}=\mathbf{X}^{(i)}\bm{g}^{(i)}_{m}+\bm{w}^{(i)}_{m}=\sum_{k\in\mathcal{K}}g^{(i)}_{km}\bm{x}^{(i)}_{k}+\bm{w}^{(i)}_{m}$, where $\mathbf{X}^{(i)}\hskip-2.84526pt=[\bm{x}^{(i)}_{1}\bm{x}^{(i)}_{2}\cdots\bm{x}^{(i)}_{K}$] is the $T_{p}\times K$ pilot matrix and $\bm{g}^{(i)}_{m}\hskip-2.84526pt=[g^{(i)}_{1m}\;g^{(i)}_{2m}\cdots g^{(i)}_{Km}$$^{\top}$ is the channel vector (of length $K$) for O-RU $m$. Here, $\bm{w}^{(i)}_{m}\sim\mathcal{CN}(\mathbf{0},\sigma^{2}\mathbf{I}_{T_{p}})$ is the zero-mean complex Gaussian noise vector of length $T_{p}$ with covariance $\sigma^{2}\mathbf{I}_{T_{p}}$, where $\mathbf{I}_{n}$ is the $n\times n$ identity matrix.
@@
$\hskip-5.69054pt\widehat{g}^{(i)}_{km}\hskip-2.84526pt=\hskip-1.42262pt(\hskip-1.42262pt\bm{x}_{k}^{(i)})^{\text{H}}\bm{y}^{(i)}_{m}\hskip-1.42262pt=\hskip-4.2679pt\sum_{k^{\prime}\in\mathcal{K}}g^{(i)}_{k^{\prime}m}(\hskip-1.42262pt\bm{x}_{k}^{(i)})^{\text{H}}\bm{x}^{(i)}_{k^{\prime}}\hskip-1.42262pt+\hskip-1.42262pt(\hskip-1.42262pt\bm{x}_{k}^{(i)})^{\text{H}}\bm{w}^{(i)}_{m}\hskip-1.42262pt=\hskip-1.42262ptg^{(i)}_{km}\hskip-1.42262pt+\hskip-4.2679pt\sum_{\begin{subarray}{c}k^{\prime}\in\mathcal{K}\\ k^{\prime}\neq k\end{subarray}}g^{(i)}_{k^{\prime}m}(\hskip-1.42262pt\bm{x}_{k}^{(i)})^{\text{H}}\bm{x}^{(i)}_{k^{\prime}}\hskip-1.42262pt+\hskip-1.42262pt(\hskip-1.42262pt\bm{x}_{k}^{(i)})^{\text{H}}\bm{w}^{(i)}_{m}$. (4)
@@ II-D Data Transmission Model
$\hskip-5.69054pt\mathsf{SINR}^{\mathsf{u}}_{k}=\frac{\tiny\rho_{k}\left|\mathbb{E}\left[\sum_{m\in{\mathcal{M}^{\mathsf{UE}}_{k}}}(\widehat{g}^{(i)}_{km})^{*}g^{(i)}_{km}\right]\right|^{2}}{\tiny\displaystyle\sum_{k^{\prime}\in\mathcal{K}}\rho_{k^{\prime}}\mathbb{E}\Bigg[\bigg|\sum_{m\in\mathcal{M}^{\mathsf{UE}}_{k}}(\widehat{g}^{(i)}_{km})^{*}g^{(i)}_{k^{\prime}m}\bigg|^{2}\Bigg]-\rho_{k}\Bigg|\mathbb{E}\Bigg[\sum_{m\in{\mathcal{M}^{\mathsf{UE}}_{k}}}(\widehat{g}^{(i)}_{km})^{*}g^{(i)}_{km}\Bigg]\Bigg|^{2}+\sigma^{2}_{\mathsf{u}}\sum_{m\in{\mathcal{M}^{\mathsf{UE}}_{k}}}\mathbb{E}\left[|\widehat{g}^{(i)}_{km}|^{2}\right]},$ (6)
@@
[…]
@@ II-E Problem Formulation
$\mathsf{MSE}^{(i)}_{k}$ $=\mathbb{E}\bigg[\sum_{m\in\mathcal{M}^{\mathsf{UE}}_{k}}\Big|\widehat{g}^{(i)}_{km}-g^{(i)}_{km}\Big|^{2}\bigg]=\hskip-5.69054pt\sum_{m\in\mathcal{M}^{\mathsf{UE}}_{k}}\mathbb{E}\left[\Big|\widehat{g}^{(i)}_{km}-g^{(i)}_{km}\Big|^{2}\right]$
$\hskip-28.45274pt=\hskip-5.69054pt\sum_{m\in\mathcal{M}^{\mathsf{UE}}_{k}}\mathbb{E}\Bigg[\Big|\sum_{\begin{subarray}{c}k^{\prime}\in\mathcal{K}\\ k^{\prime}\neq k\end{subarray}}g^{(i)}_{k^{\prime}m}(\bm{x}_{k}^{(i)})^{\text{H}}\bm{x}^{(i)}_{k^{\prime}}+(\bm{x}_{k}^{(i)})^{\text{H}}\bm{w}^{(i)}_{m}\Big|^{2}\Bigg]=\hskip-5.69054pt\sum_{m\in\mathcal{M}^{\mathsf{UE}}_{k}}\sum_{\begin{subarray}{c}k^{\prime}\in\mathcal{K}\\ k^{\prime}\neq k\end{subarray}}\beta^{(i)}_{k^{\prime}m}\Big|(\bm{x}_{k}^{(i)})^{\text{H}}\bm{x}^{(i)}_{k^{\prime}}\Big|^{2}+\sigma^{2},$ (9)
@@
$\text{s.t.}\;\;\bm{x}^{(i)}_{k}\in\mathcal{T}^{(i)}_{u},\;\;\forall k\in\mathcal{K}^{\mathsf{DU}}_{u},$ (11)
$\hskip 17.07164pt\|\bm{\phi}^{(i)}_{u,t}\|^{2}_{2}=1,\left(\bm{\phi}_{u,t}^{(i)}\right)^{\hskip-1.99168pt\text{H}}\hskip-2.84526pt\bm{\phi}^{(i)}_{u,t^{\prime}}=0\text{ if }t\neq t^{\prime},\;\;\forall t,t^{\prime}=1,2,\ldots,T_{p}.$ (12)
@@ III-B Preliminaries on Multi-agent Deep Reinforcement Learning
MA-DRL addresses scenarios where multiple agents perform simultaneous decision-making based on a Markov game model [44]. For our decentralized PA problem, we define MA-DRL using a tuple $(\{\bm{S}^{(\ell)}_{u}\}_{u\in\mathcal{U}}$, $\{\bm{a}^{(\ell)}_{u}\}_{u\in\mathcal{U}}$, $\{r^{(\ell)}_{u}\}_{u\in\mathcal{U}})$, where $\bm{S}^{(\ell)}_{u}$, $\bm{a}^{(\ell)}_{u}$, and $r^{(\ell)}_{u}$ are respectively the state, action, and reward of the agent $u$ during the $\ell$-th near-RT loop. For each loop $\ell$, agent $u$ with a state $\bm{S}^{(\ell)}_{u}$ makes an action $\bm{a}^{(\ell)}_{u}$ to interact with the environment. Subsequently, the agent makes an observation and computes a reward $r^{(\ell)}_{u}$ which helps to find the next state $\bm{S}^{(\ell+1)}_{u}$.
[…]
@@ III-C1 States
$\bm{\Phi}_{u}^{(\ell)}]_{k,t}=\begin{cases}1&\text{if }\bm{x}^{(N_{\mathsf{n}}\ell)}_{k}=\bm{\phi}^{(N_{\mathsf{n}}\ell)}_{u,t},\\ 0&\text{otherwise}.\end{cases}$ (13)
@@ III-C3 Rewards
where $p=\widetilde{p}^{(\ell)}_{u}$ by the availability of inter-DU message. The mapping function (16) converts the observation data into a reward range such that lower values of $p$ are rewarded higher. Here $p_{\text{min}},p_{\text{max}}]$ is the range of observation data, which we assume is set by the non-RT RIC.
@@
[…]
@@
$\hskip-11.38109pt\bm{a}_{u}^{(\ell)}=\operatorname*{arg\,min}_{\bm{a}_{u}\in\mathcal{A}_{u}}\;\sum_{k\in\mathcal{K}^{\mathsf{DU}}_{u}}\sum_{m\in\mathcal{M}^{\mathsf{UE}}_{k}}\mathbb{E}[p^{(\ell)}_{km}]=\operatorname*{arg\,min}_{\bm{a}_{u}\in\mathcal{A}_{u}}\frac{1}{N_{\mathsf{n}}}\sum_{k\in\mathcal{K}^{\mathsf{DU}}_{u}}\sum_{m\in\mathcal{M}^{\mathsf{UE}}_{k}}\sum_{n=0}^{N_{\mathsf{n}}-1}\mathbb{E}\left[\left|\widehat{g}^{(N_{\mathsf{n}}\ell+n)}_{km}\right|^{2}\right],$ (21)
@@ III-D MA-DRL-based Pilot Assignment Scheme
1Input: Pilot length $T_{p}$, number of RT loops NN, number of RT loops per near-RT loop $N_{\mathsf{n}}$, number of internal loops $L$, set of users managed by O-DU $u$ $\mathcal{K}^{\mathsf{DU}}_{u}$, set of O-RUs managed by O-DU $u$ $\mathcal{M}^{\mathsf{DU}}_{u}$, set of users served by O-RU $m$ $\mathcal{K}^{\mathsf{RU}}_{m}$, set of O-RUs serving the user $k$ $\mathcal{M}^{\mathsf{UE}}_{k}$, training period, update period
@@ 13  Update the PA according to $\bm{a}^{(\ell)}_{u}$
16 if $\text{mod}(\ell,\text{training period})=0$ then generate a batch from the memory and train $\bm{\theta}^{\mathsf{tr}}_{u}$ via SGD on (24)
17  if $\text{mod}(\ell,\text{update period})=0$ then set $\bm{\theta}^{\mathsf{ta}}_{u}$ = $\bm{\theta}^{\mathsf{tr}}_{u}$